Anhydrous copper(II) sulfate is difficult to dry completely. What mass of copper(II) sulfate would remain after removing 90% of the water from 250. g of CuSO4×5H2O? 


1
Expert's answer
2022-03-13T15:18:24-0400

M(CuSO4×5H2O) = 250 g/mole;

M(CuSO4) = 160 g/mole;

X(CuSO4) = M(CuSO4)/M(CuSO4×5H2O) = 160/250 = 0.64 = 64 %;

X(H2O) = 1 - X(CuSO4) = 1 - 0.64 = 0.36 = 36 %

m(CuSO4) = msubstance * X(CuSO4) = 250 * 0.64 = 160 g;

m(H2O) = msubstance * X(H2O) = 250 * 0.36 = 90 g;

Find the water content after removing 90% of the water from 250 g of CuSO4×5H2O:

m(H2O)remain = m(H2O) * 100%-90%/100% = 90 * 0.1 = 9 g;

Find the mass of copper(II) sulfate would remain after removing 90% of the water from 250 g of CuSO4×5H2O:

m(substance)remain = m(CuSO4) + m(H2O)remain = 160 + 9 = 169 g.

Answer: 9 g.